. Gravimetric analysis is widely used in the mining and metal-alloy industries to determine

chemical composition. An iron ore is analyzed for the iron (FW = 55.845 g/mol) content by

converting it to Fe2O3 (FW = 159.69 g/mol) precipitate.

a) If the iron content is in the range of 8% to 15%, what is the minimum mass of

sample (in four significant result) required to obtained 100.0 mg of precipitate?

[4 Marks]

b) Calculate the percent by mass of iron in the precipitate if 1.52 g sample yields

0.126 g Fe2O3.

[4 Marks]

c) Explain how the von Weimarn ratio can affect the particle sizes of precipitate.

[2 Marks]


1
Expert's answer
2021-12-07T07:52:02-0500

atomic mass of Fe is 55.85 g/mol

molar mass of Fe2O3 is 159.69 g/mol

mass percentage of Fe in Fe2O3=55.85*2/159.69=69.95%

100mg Fe2O3 contains Fe 100*69.95%=69.95 mg

Fe occupies 10-14% of sample mass

sample mass in the range of 69.95/14%-69.95/10%, namely 499.6-699.5 mg

minimal sample mass is 499.6mg



According to Von Weimarn, the mean magnitude of the individual crystals of precipitates will progressively decrease with progressively increasing concentrations of the reacting solutions providing the process of direct crystallization is complete.

Therefore, the higher the concentration of barium sulfate solution, the lower the mean magnitude of barium sulfate crystals.